OpenStudy (anonymous):

\[\int\limits_{}^{} \cos^2 \theta \sin \theta \]

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

\[\int\limits_{}^{} \cos^2 \theta \sin \theta \]

OpenStudy (anonymous):

\int\limits_{}^{} \cos^2 \theta \sin \theta d theta

OpenStudy (anonymous):

\[d \theta \] at the end